OpenStudy (cwtan):

Please help me to prove (1-sin x)/(1+sin x)= tan^2 (pi/4 - x/2).

OpenStudy (cwtan):

sad... nobody helps me.....

OpenStudy (anonymous):

Wait, I will help you. Needs a lot of typing.

OpenStudy (cwtan):

I almost want to close this question...

OpenStudy (anonymous):

\[\frac{1-\sin (x)}{\sin (x)+1}=\frac{(1-\sin (x)) (\sin (x)+1)}{(\sin (x)+1) (\sin (x)+1)}=\frac{1-\sin ^2(x)}{(\sin (x)+1)^2}= \]

OpenStudy (anonymous):

\[\frac{1-\sin ^2(x)}{(\sin (x)+1)^2}= \frac {\cos^2(x)}{(\sin (x)+1)^2} \]

OpenStudy (anonymous):

We need to show that now \[ \frac{\cos (x)}{\sin (x)+1}=\tan \left(\frac{\pi }{4}-\frac{x}{2}\right) \]

OpenStudy (anonymous):

\[\tan \left(\frac{\pi }{4}-\frac{x}{2}\right)=\frac{\sin \left(\frac{\pi }{4}-\frac{x}{2}\right)}{\cos \left(\frac{\pi }{4}-\frac{x}{2}\right)}=\frac{\cos \left(\frac{x}{2}\right)-\sin \left(\frac{x}{2}\right)}{\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)}=\\ \frac{\left(\cos \left(\frac{x}{2}\right)-\sin \left(\frac{x}{2}\right)\right) \left(\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)\right)}{\left(\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)\right) \left(\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)\right)}=\\ \frac{\cos^2\left(\frac{x}{2}\right)-\ \sin^2\left(\frac{x}{2}\right)}{ \left(\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)\right)^2}= \frac{\cos (x)}{\sin ^2\left(\frac{x}{2}\right)+\ cos ^2\left(\frac{x}{2}\right)+2 \sin \left(\frac{x}{2}\right) \cos \left(\frac{x}{2}\right)}=\\ \frac{\cos (x)}{1+2 \sin \left(\frac{x}{2}\right) \cos \left(\frac{x}{2}\right)}=\frac{\cos (x)}{1+\sin(x)} \]

OpenStudy (anonymous):

We are done.

OpenStudy (cwtan):

Wow ~ Thanks!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Another quicker way \[ \tan \left(\frac{\pi }{4}-\frac{x}{2}\right)=\frac{1-\tan \left(\frac{x}{2}\right)}{\tan \left(\frac{x}{2}\right)+1}=\frac{\cos \left(\frac{x}{2}\right)-\sin \left(\frac{x}{2}\right)}{\sin \left(\frac{x}{2}\right)+\cos \left(\frac{x}{2}\right)}=\frac{\cos (x)}{\sin (x)+1} \]